OpenStudy (anonymous):

Solve for x: 2/3(x-4)=2x

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }(x-4)=2x\]

OpenStudy (anonymous):

@phi @thomaster @Luigi0210 help please?

OpenStudy (campbell_st):

multiply both sides of the equation by 3 and you'll have 2(x - 4) = 6x you can either divide both sides by 2... or distribute then collect like terms and solve for x

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

So it would be x=-2?

OpenStudy (campbell_st):

yes... that's the solution I had